Which equation does not represent a line of symmetry for the circle with equation x^2+y^2=100?
a. x = 0
b. y= 1/2 x
c. y = x
d. y = x + 1

Answer by ikleyn(52781)

.
Which equation does not represent a line of symmetry for the circle with equation x^2+y^2=100?

a. x = 0

b. y= 1/2 x

c. y = x

d. y = x + 1                  <<<---=== This one.